# White Men Lead, Black Women Help: Uncovering Gender, Racial, and Intersectional Bias in Language Agency

### v1

@@ 4.2.1 Datasets
Personal Biographies   To study language agency gender biases, we use Bias in Bios De-Arteaga et al. (2019), a biography dataset extracted from Wikipedia pages. Since the biography data for different professions are significantly imbalanced, we randomly sample $120$ biographies for each gender for each of the professions. A full list of professions in the pre-processed dataset is in Appendix C, Table 13. Since the dataset of Wikipedia-based human-written biographies does not include racial information, we follow the Context-Less Generation (CLG) setting in  Wan et al. (2023) and utilize ChatGPT to generate a biographies dataset for males and females of $4$ racial groups: White, Black, Hispanic, and Asian. We then use this dataset to study gender, racial, and intersectional language agency biases in LLM-generated biographies.
@@
Professor Reviews   We investigate language agency bias in human-written professor reviews using an open-access sample dataset of student-written reviews for professors  [^2] https://github.com/x-zhe/RateMyProfessor_Sample_Dataset, which was web-crawled from the RateMyProfessor website  [^3] https://www.ratemyprofessors.com/. We begin preprocessing by removing the majority of data entries without professors’ gender information. Since the remaining data is scarce and unevenly distributed across genders and departments, we remove data from departments with less than $10$ reviews for male or female professors. No further re-sampling or balancing is conducted on the remaining dataset, since that would remove most data entries. A full list of departments and corresponding gender distributions of professor reviews in the pre-processed dataset is provided in Appendix C, Table 14. Since the RateMyProfessor dataset does not include racial information, we again used ChatGPT under the CLG setting to generate a professor review dataset with synthetic comments for male and female professors of the $4$ above-mentioned racial groups, and from the same set of departments as in the RateMyProfessor data. Then, we use the dataset to comprehensively study language agency biases in model-synthesized professor reviews.
Reference Letters   We analyze language agency gender bias using the ChatGPT-generated reference letter dataset in  Wan et al. (2023)’s work. Since  Wan et al. (2023)’s dataset does not include racial information, we again adopted the CLG approach and employed ChatGPT to generate a reference letter dataset for individuals, with both gender and racial specifications. We then used this synthetic dataset to investigate language agency bias in gender, race, and intersectional identities in LLM-generated letters of recommendation.
4.2.2 Reported Metrics
For each dataset for experiments, we first calculate the average percentage of agentic and communal sentences in texts describing different social groups. We then mainly report the average ratio gap between agentic and communal sentences, to better reflect the overall language agency level. For gender bias evaluation, we further conducted statistical t-testing with alternative hypothesis ($H_{1}$) being that male texts are higher in agency level than female texts, and null hypothesis ($H_{0}$) being no significant agency differences. Statistical testing results are in Appendix D, Table 18.
4.3 Results
Below, we illustrate experiment results revealing different aspects of language agency biases. All results are averaged on random seeds 00, $1$, and $2$.
4.3.1 Gender Bias
Table 2 shows results of gender biason the $6$ human-written and LLM-generated datasets. Across all datasets, texts describing males are remarkably higher in language agency level than those describing females. Additionally, language agency gender biases found on human-written texts in our study align with findings of social science studies, showing that language agency is effective in capturing biases in implicit language styles. Furthermore, both ratio-based (Table 2) and statistical testing results (Appendix D Table 18) show that LLM-generated texts demonstrate more severe language agency gender biases than human-written texts. To better visualize overall results on the $3$ LLM-generated datasets, we use $3$ Kernel Density Estimate graphs in the upper row of Figure 2. We observe that texts about females possess higher density to the lower language agency end (i.e. left), whereas texts about males have higher density to the higher agency level end (i.e. right).
Dataset Gender AvgAvgAvg
Agen. Comm. Gap
Biography (Human) Female $63.81$ $36.19$ $27.61$
Male 68.87 31.13 37.73
Biography (LLM) Female $38.28$ $61.72$ $-23.45$
Male 42.52 57.48 -14.96
Professor Review (Human) Female $77.84$ $22.16$ $55.67$
Male 78.76 21.24 57.53
Professor Review (LLM) Female $32.79$ $67.21$ $-34.42$
Male 36.07 63.93 -27.85
Reference Letter (Wan et al.) Female $55.15$ $44.85$ $10.30$
Male 57.47 42.53 14.94
Reference Letter (LLM) Female $53.26$ $46.74$ $6.52$
Male 57.92 42.08 15.85
Table 2: Experiment results for language agency gender bias in natural language texts. Results with the highest language agency level for each dataset are in bold. Overall, texts describing males demonstrate a higher level of language agency than texts describing females.
Bias in Biography   We investigate gender biases in both human-written and LLM-generated biographies, and observe that the average ratio gap between agentic and communal sentences in male biographies is higher than that of females. Our observations align with findings in prior social science work (Wagner et al., 2016), which revealed gender bias towards females in Wikipedia biographies. We further stratify analysis on the human-written Bias in Bios dataset based on the professions of individuals in biographies in Appendix D Table 20, and visualize the top $8$ most biased occupations as overlap horizontal bar graphs in Figure 1. Drastic language agency gender bias are found for pastor, architect, and software engineer. Interestingly, real-world reports have also demonstrated male dominance and gender bias in these occupations (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Observations on language agency gender bias in biographies align with social biases, showing the effectiveness of the proposed language agency approach in capturing biases in implicit language traits.
Bias in Professor Review   We study gender biases in both human-written and LLM-generated reviews for professors. On both human- and LLM-written professor reviews, we discover that reviews for male faculties possess higher agency than those for female faculties, aligning with gender bias observations on real-world professor ratings (Schmidt, ; Roper, 2019; Macnell et al., 2014). We further stratify our analysis on the human-written RateMyProfessor dataset based on the professors’ departments in Appendix D Table 19, and visualize the top $8$ most biased departments in Figure 1. The greatest language agency biases are observed in reviews for professors in departments such as Accounting, Sociology, and Chemistry; all $3$ departments have been proven to be male-dominated (200, 2009; Girgus, ; Seijo, ). Again, gender bias in language agency observed in our study aligns with real-world biases, further validating the effectiveness of language agency in capturing biases.
Bias in Reference Letter   On both Wan et al. (2023)’s dataset and our LLM-generated reference letter dataset, we observe that the average ratio gap between agentic and communal sentences in letters for male candidates is higher than that for female candidates. This observation aligns with prior works’(Madera et al., 2009, 2019; Khan et al., 2021; Wan et al., 2023) conclusions, pointing towards significant language agency gender biases towards females in letters of recommendation.
Dataset Race AvgAvgAvg
Agen. Comm. Gap
Biography (LLM) White $41.81$ $58.19$ $-16.38$
Black $36.39$ $63.61$ $-27.22$
Hispanic $39.06$ $60.94$ $-21.88$
Asian 44.33 55.67 -11.34
Professor Review (LLM) White $34.62$ $65.38$ $-30.76$
Black $31.35$ $68.65$ $-37.30$
Hispanic $35.84$ $64.16$ $-28.33$
Asian 35.92 64.08 -28.16
Reference Letter (LLM) White 57.50 42.50 15.00
Black $54.54$ $45.46$ $9.08$
Hispanic $55.31$ $44.69$ $10.63$
Asian $55.01$ $44.99$ $10.03$
Table 3: Experiment results for language agency racial bias in natural language texts. Highest language agency level for each dataset is in bold. Texts describing Black individuals have the lowest language agency level compared with other racial groups.
Figure 2: Kernel Density Estimate (KDE) plot of the average ratio gap between agentic and communal sentences on the 6 investigated datasets. The 3 figures in the upper row illustrate language agency gender bias, in which texts describing males have greater density towards the higher language agency end (i.e. to the right). The 3 figures in the lower row illustrate language agency racial bias, in which texts describing Black individuals tend to have greater density towards the lower language agency end (i.e. to the left).
4.3.2 Racial Bias
Table  3 shows experiment results for racial biases in language agency across $3$ LLM-generated datasets with racial information in our study. Overall, we observe that (1) reference letters possess a larger proportion of agentic sentences, whereas biographies and professor reviews mostly contain more communal descriptions, and (2) on all investigated datasets, language describing Black individuals has the lowest level of agency than that for other racial groups. We visualize experiment results on each dataset using $3$ Kernel Density Estimate graphs in the lower row of Figure 2. Across datasets, texts about Black individuals have a higher density than other racial groups in the lower language agency end (i.e. left), aligning with observations in quantitative results.
Bias in Biography   On the LLM-generated biography dataset, biographies of Asian and White individuals demonstrate the highest level of language agency. Shockingly, there is a $15.88$ difference between the agentic-communal ratio gaps for Asian biographies and Black biographies, indicating a strong racial bias towards Black individuals.
Bias in Professor Review   Results on the LLM-synthesized professor review dataset also show biases towards Black faculties. A $9.14$ discrepancy is observed between agentic-communal ratio gaps of reviews for Asian professors and for Black professors. Interestingly, studies on real-world professor ratings also found that Black professors received more negative reviews from students (Reid, 2010).
Bias in Reference Letter   We observe that highest language agency levels are observed in LLM-generated reference letters for White individuals, whereas Black individuals receive letters with the lowest language agency level. Agentic-communal ratio gap of reference letters for White individuals is $5.92$ higher than Black individuals, $4.37$ higher than Hispanic individuals, and $4.97$ higher than Asian individuals. Such racial biases in LLM-written letters of recommendation also aligns with previous social science findings (Powers et al., 2020; Chapman et al., 2022).
4.3.3 Intersectional Bias
Different from prior works, which mostly focused on single bias aspects, we are amongst the first to investigate intersectional biases in LLM-generated texts through the lens of language agency. Specifically, we use the $3$ LLM-generated datasets—which contains both gender and racial information—to study intersectional gender and racial biases. Results show that texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. To visualize results, we plot the average ratio gap between agentic and communal sentences in texts describing different genders for different racial groups as overlap horizontal bar graphs in Figure 1. Both quantitative results and visualized illustrations show intersectional racial and gender biases across 3 datasets.
@@ 4.3.3 Intersectional Bias
Figure 3: Visualization of the average ratio gap between agentic and communal sentences in the 3 datasets for different intersectional genders and racial groups. Agency levels of texts describing males are notably higher than those describing females in each racial group.
Dataset Race Gender AvgAvgAvg
Agen. Comm. Gap
Biographies (LLM) White Male $43.87$ $56.13$ $-12.27$
Female $39.76$ $60.24$ $-20.49$
Black Male $37.23$ $62.77$ $-25.53$
Female $35.55$ $64.45$ $-28.91$
Hispanic Male $42.30$ $57.70$ $-15.40$
Female $35.82$ $64.18$ $-28.36$
Asian Male 46.68 53.32 -6.64
Female $41.98$ $58.02$ $-16.03$
Table 4: Experiment results for intersectional gender and racial bias in language agency in biographies. Highest language agency level for each dataset is in bold. Biographies for Black females are lowest in language agency level, whereas biographies for Asian males are highest in agency.
Bias in Biographies   Table 4 shows results for intersectional biases in LLM-generated biographies. We observe that (1) biographies of Asian males demonstrate the highest level of language agency, (2) biographies of Black females show the lowest level of agency, and (3) within each racial group, language agency is notably higher in male biographies than in female ones. Similar findings have also been made in social science works, unveiling intersectional racial and gender biases in biographies (Field et al., 2022; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015)
Professor Review   Results for intersectional biases in LLM-synthesized professor reviews are shown in Table 5 While reviews for Hispanic male professors have the highest language agency, reviews for Black female professors again show the lowest agency level. This aligns with findings of prior social science studies, which showed that female professors from racial minority groups tend to receive lower scores in student ratings (Chávez and Mitchell, 2020). Additionally, within racial groups, reviews for female professors have lower levels of language agency than male faculties.
Dataset Race Gender AvgAvgAvg
Agen. Comm. Gap
Professor Review (LLM) White Male $36.01$ $63.99$ $-27.98$
Female $33.23$ $66.77$ $-33.54$
Black Male $31.95$ $68.05$ $-36.10$
Female $30.74$ $69.26$ $-38.51$
Hispanic Male 38.52 61.48 -22.95
Female $33.15$ $66.85$ $-33.71$
Asian Male $37.81$ $62.19$ $-24.39$
Female $34.03$ $65.97$ $-31.93$
Table 5: Experiment results for intersectional gender and racial bias in language agency in professor reviews. Highest language agency level for each dataset is in bold. Reviews for Black female professors demonstrate the lowest language agency, while reviews for Hispanic male faculties have the highest agency level.
Dataset Race Gender Avg % Avg % Avg
Agen. Comm. Gap
Reference Letter (LLM) White Male 59.88 40.12 19.75
Female $55.12$ $44.88$ $10.24$
Black Male $56.74$ $43.26$ $13.49$
Female $52.34$ $47.66$ $4.67$
Hispanic Male $57.64$ $42.36$ $15.27$
Female $52.99$ $47.01$ $5.98$
Asian Male $57.44$ $42.56$ $14.87$
Female $52.59$ $47.41$ $5.19$
Table 6: Experiment results for intersectional gender and racial bias in language agency in reference letters. Highest language agency level for each dataset is in bold. Recommendation letters for Black females again are the lowest in language agency; letters for White males are the highest in agency level.
Reference Letters   Table 6 shows results for intersectional biases in LLM-generated reference letters. Results show that (1) reference letters for white males have the highest levels of language agency, (2) letters for Black females again demonstrate the lowest level of agency, and (3) within the same racial group, level of language agency is remarkably lower in letters of recommendation for females than those for males.

### v3

@@ 4.1.1 Human-Written Texts: Dataset Details
Personal Biographies   We use Bias in Bios De-Arteaga et al. (2019), a biography dataset extracted from Wikipedia pages. Since the biography data for different professions are significantly imbalanced, we randomly sample $120$ biographies for each gender for each of the professions. A full list of professions in the pre-processed dataset is in Appendix D, Table 12.
@@
Professor Reviews   We use an open-access sample dataset of student-written reviews for professors  [^1] https://github.com/x-zhe/RateMyProfessor_Sample_Dataset, which was web-crawled from the RateMyProfessor website  [^2] https://www.ratemyprofessors.com/. We first remove the majority of data entries without professors’ gender information. Since the remaining data is scarce and unevenly distributed across genders and departments, we remove data from departments with less than $10$ reviews for either gender. A full list of departments and corresponding gender distributions of professor reviews in the pre-processed dataset is provided in Appendix D, Table 13.
Reference Letters   Since we were not able to find publicly available human-written reference letter datasets, we choose to use the reference letter dataset from the Context-Based Generation (CBG) setting in  Wan et al. (2023)’s work. The CBG setting provides a paragraph of biographical information about individuals (e.g. career, life) to prompt LLMs for letter generations, which is very similar to real-world reference-letter-writing scenarios. Therefore, we use  Wan et al. (2023)’s dataset as a proxy for human-written reference letters.
4.1.2 Comparison Results
Table 2 and Figure 2 show language agency gender biases in human-written and LLM-generated biographies, professor reviews, and reference letters. We report the gender differences (Male - Female) in the intra-group agency-communal ratio gaps. Below are our observations:
Gender biases persist in language agency levels in both human-written and LLM-generated texts. Across all categories of texts, languages describing males are remarkably higher in language agency level than those describing females.
Biases observed in human-written texts in our study align with findings of social science studies. We stratify analysis on the human-written biography dataset based on professions in Appendix E, and found that occupations with greatest biases—such as pastor, architect, and software engineer—are also reported by real-world studies to be male-dominated (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Academic departments in which the highest language agency biases in professor reviews are identified—such as Accounting, Sociology, and Chemistry—have also been proven for male dominance (200, 2009; Girgus, ; Seijo, ). The alignment between our observations and real-world inequalities further validates the effectiveness of language agency in capturing social biases.
@@ 4.1.2 Comparison Results
Figure 3: Visualization of the average ratio gap between agentic and communal sentences in the 3 datasets for different intersectional genders and racial groups.
LLM-generated texts demonstrate more severe language agency gender biases than humans. As shown in Figure  2 and Table 2, for all 3 text categories, the highest gender bias levels, as measured by the gender differences in intra-group ratio gaps between agentic and communal sentences, are observed in LLMs. For professor reviews and reference letters, human-written texts demonstrate remarkably less bias than all 3 LLMs investigated. This warns of the potential propagation and even amplification of social biases in LLM-generated texts.
Model Text Type Bias Dimension
Gender$(\downarrow 0)$ Race$(\downarrow 0)$ Intersectional$(\downarrow 0)$ Overall $(\downarrow 0)$
ChatGPT Biography 38.06 47.79 66.31 50.72
  + Mitigation 29.55 14.09 45.94 29.86 (-20.86) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-20.86)}}}$
Professor Review 22.25 19.35 32.14 24.58
  + Mitigation 15.50 34.90 62.26 37.55 (+12.97) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+12.97)}}}$
Reference Letter 43.56 8.02 32.16 27.91
  + Mitigation 3.15 51.36 62.79 39.10 (+11.19) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+11.19)}}}$
Average 34.62 25.05 43.54 34.40
  + Mitigation 16.07 33.45 57.00 35.50 (+1.10) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+1.10)}}}$
Mistral Biography 60.29 29.99 61.36 50.55
  + Mitigation 19.02 22.9 32.71 24.88 (-25.67) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-25.67)}}}$
Professor Review 36.61 48.33 63.14 49.36
  + Mitigation 122.53 16.49 99.19 79.40 (+30.04) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+30.04)}}}$
Reference Letter 59.06 7.90 45.63 37.53
  + Mitigation 69.88 47.24 83.62 66.91 (+29.38) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+29.38)}}}$
Average 51.99 28.74 56.71 45.81
  + Mitigation 70.48 28.88 71.84 57.06 (+11.25) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+11.25)}}}$
Llama3 Biography 37.10 26.82 47.40 37.11
  + Mitigation 34.67 58.67 83.37 58.90 (+12.79) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+12.79)}}}$
Professor Review 68.31 85.51 125.00 92.94
  + Mitigation 10.79 9.3 22.09 14.06 (-78.88) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-78.88)}}}$
Reference Letter 44.93 26.29 49.94 40.39
  + Mitigation 23.65 20.3 43.37 29.11 (-11.28) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-11.28)}}}$
Average 50.11 46.20 74.11 56.81*
  + Mitigation 23.04 29.42 49.61* 34.02 (-22.79) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-22.79)}}}$
Table 3: Experiment results for gender, racial, and intersectional bias in language agency of 3 investigated LLMs, across 3 text generation tasks. Highest bias level for each task for each LLM is underlined. Overall bias level across all tasks and all bias dimensions for each LLM is in bold. Llama3 demonstrates the highest overall bias in language agency (*).
4.2 Findings 2: LLMs Suffer From Gender, Racial, and Especially Intersectional Biases in Language Agency
Table  3 demonstrates full results for gender, racial, and intersectional biases in language agency for biographies, professor reviews, and reference letters generated by the investigated 3 LLMs. We also visualize the average agentic-communal ratio gap in texts describing different gender and racial intersectional groups as overlapping horizontal bar graphs in Figure 4.
In the gender bias dimension, LLMs tend to depict males with more agentic language than females. As discussed in Section  4.1, all 3 LLMs possess notable levels of gender differences in agentic-communal ratio gaps. Table  3 further shows high variances of agency levels across gender groups. Both observations reveal notable language agency gender biases in LLM-generated texts.
In the racial bias dimension, LLM-generated texts for colored individuals are often remarkably less agentic than those for White individuals. Across all generation tasks, LLM-written texts about colored individuals have notably lower agency level than those for White individuals. For instance, as shown in Figure  3, Black professors receive reviews with the lowest agency levels in Chatgpt- and Llama3-generated reviews; huge discrepancies can be observed between agentic-communal ratio gaps in reviews for Black faculties and for professors of other races. Interestingly, studies on real-world professor ratings also found that Black professors received more negative reviews from students (Reid, 2010). Similarly, LLM-generated reference letters for White individuals are highest in agency, whereas Black individuals receive letters with the lowest language agency level, aligning with previous social science findings on racial biases (Powers et al., 2020; Chapman et al., 2022).
In intersectional bias dimension, texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. Both quantitative results in Appendix E Tables 20, 22,24 and visualized illustrations in Figure  3 show severe intersectional biases across all LLMs on all generation tasks—those who are at the intersection of gender and racial minority groups are the most vulnerable to biases in language agency. For instance, ChatGPT- and Llama3-generated reviews for Black female professors show the lowest level of agency across all intersectional groups. Interestingly, we observe that on all text generation tasks, language agency is notably higher in texts about males within each racial group (e.g. Black males are described with more agentic language than Black females). These observations further align with prior social science findings on intersectional biases targeting gender and racial minority groups in texts (10.1145/3485447.3512134; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015; Chávez and Mitchell, 2020).
4.3 Findings 3: All LLMs Show Agency Biases, and Llama3 is the Most Biased One
As shown by the overall quantitative results (bottom right cell in rows for each model) of biases in each dimension, for each text generation task, and for all 3 models, Llama3 suffers from the most remarkable language agency biases. Through inspecting individual bias categories, it is shocking to see that Llama3 in fact carries the highest level of biases in 2 of the 3 investigated text generation tasks—biography and reference letter—across gender, race, and intersectional dimensions. This warns that although recently-developed LLMs all demonstrate fascinating text generation abilities, they can demonstrate drastically different levels of fairness issues. Using such technology without scrutiny can result in the propagation of severe social harm.
Model Text Type Problem Type Race Gender Diff. (M-F) Gender Diff. (post-mitigate)
ChatGPT Biography Amplification Black 3.38 18.68
Professor Review Overshooting White 5.56 -16.84
Black 2.41 -9.91
Mistral Professor Review Amplification White 7.02 14.59
Black 8.02 17.93
Asian 7.99 23.00
Reference Letter Amplification Black 4.42 12.95
Asian 10.87 15.92
Llama3 Biography Amplification White 6.06 10.59
Black 6.21 11.24
Table 4: Selected quantitative results showing problems with bias amplification and bias overshooting for prompt-based mitigation methods in LLMs. For each racial group, we report the gender difference (male-female) in agency-communal ratio gap in texts. ChatGPT demonstrates both amplification and overshooting, whereas Mistral and Llama mainly suffer from amplified language agency bias levels.
4.4 Findings 4: Simple Prompt-Based Mitigation Could Worsen Language Agency Bias.
A large body of recent research has explored the use of “ethical intervention”, or prompt-based mitigation, to resolve fairness issues in textual and multimodal generative models (Bansal et al., 2022; Ganguli et al., 2023; Huang et al., 2024; Wan and Chang, 2024). Following the previous studies, we experimented with a prompt-based bias mitigation method by appending a “fairness instruction” at the end of each text generation prompt: “When generating the {text type}, ensure that you display no biases in language agency across gender or race.”
Quantitative results in Table 3 show that prompt-based methods fail to effectively resolve language agency bias. More shockingly, we reveal that fairness instructions in prompts could even result in higher bias levels in LLM-generated texts. We found 2 main problems with prompt-based mitigation methods in LLMs: (1) bias amplification, for which even more severe biases are observed, and (2) bias overshooting, where biases are shifted towards anti-stereotype directions (e.g. biased towards males). In Table  4, we provide quantitative examples for these problems for the 3 LLMs.
These surprising findings unveil the severe drawbacks of prompt engineering as a bias mitigation method—LLMs might fail to understand what is “fair” in language agency, therefore worsening existing biases or resulting in biases in the anti-stereotypical direction.